OpenStudy (anonymous):

Justify if completing the square is a good method for solving when the Discriminant is negative.

OpenStudy (anonymous):

why not? it is as good a method as any

OpenStudy (anonymous):

is this a joke?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

no it is not a joke

OpenStudy (anonymous):

if i saw \[x^2+4x+10=0\] and didn't mind complex solutions i would write \[x^2+4x=-10\\ (x+2)^2=-10+4=-6\] and get \[x+2=\pm\sqrt{6}i\] right away